OpenStudy (anonymous):

Find an equation of the parabola with directrix y=6 and focus (0,2).

OpenStudy (amistre64):

vertex is halfway between d and f

OpenStudy (anonymous):

so directrix would be same length between vertex right?

OpenStudy (amistre64):

yeah; -4ay = x^2 is what we wanna transform; and a = half the distance between d and f 6-2 = 4/2 = 2.. a =2

OpenStudy (anonymous):

??? lol i'm lost.

OpenStudy (amistre64):

-4(2)(y-4) = x^2 accounts for the vertex i think

OpenStudy (amistre64):

the vertex is at (0,4) whic is half way between the directix and the focus

OpenStudy (amistre64):

the distance from the focus to the vertex is defined in most cases as 'a' -4(2)(y-4) = x^2 y-4 = (-1/8)x^2 y = (-1/8)x^2 +4

OpenStudy (anonymous):

ok i get everything else, just why did u multiply -4(2)?

OpenStudy (amistre64):

the focus is important; and it has to be placed in the right spot in order for the parabola to be defined appropriately; the distance between focus and vertex EQUALS the distance between vertex and directix; on here that distance = 2

OpenStudy (amistre64):

it is the geometric definition for a parabola

OpenStudy (amistre64):

4ay = x^2 is the generic version of it

OpenStudy (amistre64):

since the focus is under the vertex, it opens down; so I need a -4ay to = x^2

OpenStudy (amistre64):

since the vertex is not at the origin; we make note of it in the equation; -4a(y-4) = x^2

OpenStudy (anonymous):

we didn't learn that formula, i'm trying to get the answer off of the equation y=a(x-h)^2+k

OpenStudy (amistre64):

there are many parabolas with a vertex as 0,4

OpenStudy (amistre64):

without any other points to compart with, you have to use the the geometric definition

OpenStudy (amistre64):

all points along the curve are equidistance from a perpendicular line from the directix and the focus.

OpenStudy (anonymous):

ohhhh i see that, so vertex =0,4. and so distance is 2. so u got y-4=(-1/8)x?

OpenStudy (amistre64):

well; x^2 but yeah

OpenStudy (anonymous):

ok i kind of see that.